Manuel Kauers Hagenberg, July 25, 2003 RISC-Linz

Pronountiation

“eu”

In German, the letter combinations “eu” and “äu” are pronounced as [oı], as in English oil, toy, foil Challenge: Heute läuft der freundliche Käufer neun deutscher Häuser zu neuen Leuten.

Some words with “eu”

abscheulich – ugly bereuen – to regret die Beute – beehive deutsch – german Deutschland – Germany das Efeu – ivy die Eule – owl Euklid – ancient mathematician Euphrat – river in Iraq der Euro – european currency das Feuer – fire die Freude – joy der Freund – friend freundlich – friendly heute – today die Leute – people neu – new neulich – recently neun – nine der Räuber – robber die Reue – sorrow, contrition der Säugling – baby die Säure – acid die Scheu – awe, timidity scheuen – to avoid die Schleuse – watergate das Steuer – steering wheel die Steuer – tax teuer – expensive der Teufel – devil treu – loyal, faithfull . . . . . .

Furthermore, “eu” is frequent in

• plural forms (das Haus – die Häuser, der Kauf – die Käufe, die Haut – die Häute, …)

• some konjugated verb forms (laufen – er läuft, …)

• the pronouns euer, euch

Exercise

• Complete the list of “eu”-words from the vocabulary you know.

• Build challange sentences with as many “eu”s in them as possible.